In a galvanometer there is a deflection of 10 divisions per mA. The internal resistance of the galvanometer is 78 ohm. If a shunt of 2 ohm is connected to the galvanometer and there are 75 divisions in all on the scale of the galvanometer,calculate the maximum current which the galvanometer can read. 


Grade:9

1 Answers

Samiksha
5 years ago
We will use the formula for calculating Imax of a Galvanometer, Ig = number of divisions/unit division
Ig = \frac{75}{10}, and, Imax = \frac{(Rg+Rs) Ig}{Rs} = \frac{(78+2) 75}{(78) 10} = \frac{100}{3}\boldsymbol{A}
Imax = \frac{(Rg+Rs) Ig}{Rs} = \frac{(78+2) 75}{(78) 10} = \frac{100}{3}\boldsymbol{A}